Give exact and approximate solution to three decimal places for (x-6)^2=64
Answer by Alan3354(69443)

Give exact and approximate solution to three decimal places for (x-6)^2=64
-------
(x-6)^2=64
x-6 = 8
x = 14
-------
x-6 = -8
x = -2